Doggykid:

What is the value of r? 1.2r−1+1.3r=40

Doggykid:

@Trillboij

Doggykid:

help me bro plsz!!!!!!!!!!

alones:

Add \(\ 1\) from left side of equal sign to right side making it \(\ 1.2r+1.3r=41\)

alones:

Then\(\ 1.2+1.3= 2.5r=41\)

alones:

\[ \frac{41}{2.5}=16.4 \]

alones:

So your answer would be \[\ 16.4\]

CloroxBleach:

^^

MsBrains:

gg

umer:

16.4